A man can travel 120 miles in either of two ways. He can travel at a constant rate of 40 miles per hour, or he can travel halfway at 50 miles per hour, then slow doen to 30 miles per hour for the second 60 miles. Which way is faster, and by how much?
A) The constant rate is faster by 10 minutes or more.
B) the constant rate is faster by less than 10 minutes.
C) The two ways take exactly the same time.
D) The constant rate is slower by less than 10 minutes.
E) The constant rate is slower by 10 minutes or more.

Two cities are 800 miles apart. At 3:00pm, Plane A leaves one city,traveling toward the other city at a speed of 600 miles per hour. At 4:00 the same afternoon, Plane B leaves the first city, traveling in the same direction at a rate of 800 miles per hour. Which of the following answers represents the actual rate?
A) Plane A arrives first by an hour or more.
B) Plane A arrives first by less than an hour.
C) The two planes arrive at exactly the same time.
D) Plane A arrives after Plane B by less than an hour.
E) Plane A arrives after Plane B by an hour or more.
C)

Answer by stanbon(75887)

A man can travel 120 miles in either of two ways. He can travel at a constant rate of 40 miles per hour, or he can travel halfway at 50 miles per hour, then slow doen to 30 miles per hour for the second 60 miles. Which way is faster, and by how much?
----
Constant speed time = 120/40 = 3 hrs
Change of speed time = 60/50 + 60/30 = (6/5)+ 2 = 3 1/5 hrs = 3 hrs 12 min
------------
E) The constant rate is slower by 10 minutes or more.
============================================================
Two cities are 800 miles apart. At 3:00pm, Plane A leaves one city,traveling toward the other city at a speed of 600 miles per hour. At 4:00 the same afternoon, Plane B leaves the first city, traveling in the same direction at a rate of 800 miles per hour. Which of the following answers represents the actual rate?
----
Plane A DATA:
rate = 600 mph ; distance = 800 miles ; tme = 8/6 = 4/3 hours = 1hr 20 min
Arrival time = 3:00+ 1hr 20min = 4:20 PM
---
Plane B DATA:
rate = 800 mph ; distance = 800 miles ; time = 1 hr
Arrival time = 4:00 + 1hr = 5:00 PM
--------------------------------------
B) Plane A arrives first by less than an hour.
=============
Cheers,
Stan H.
=============